OpenStudy (anonymous):

Find the new balance, assuming that the bank charges 2 ½ % per month on the unpaid balance. Previous Balance Payment New Purchases $100 $20 $75

OpenStudy (dumbcow):

100-20 = 80 unpaid balance is 80 80*(1.025) = 82 82+75 = 157 new balance is 157